OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet. Show your work

OpenStudy (anonymous):

you need to subtract $6 from $54, then divide $8 into $48, and $48 divided by $8 is 6

OpenStudy (anonymous):

Put it in an equation please lol

OpenStudy (anonymous):

set x is the price of bracelet so 8x+6=54 8x=54-6 8x=48 x=48/8 x=6 the bracelet cost $6

OpenStudy (anonymous):

Thanks guys!!

OpenStudy (anonymous):

8c + 6 = 54